OpenStudy (anonymous):

Formula for this sequence, and explain your answer. Likewise, does this converge or diverge if we take the sum from 1 to infinity? If it converges, what does it converge to? { 1, -1/2, -1/3, 1/4, 1/5, -1/6, -1/7, 1/8, 1/9, -1/10, -1/11, 1/12, 1/13...}

OpenStudy (anonymous):

@DollyAcquah

OpenStudy (anonymous):

sorry i really this type of math :(

OpenStudy (anonymous):

That's ok :)

OpenStudy (anonymous):

I know the formula for alternating fractions, but not the one where it alternates every 2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

I believe that taking the sum from 1 to infinity instead of 0 to infinity might mean the formula is something/n. I know that (-1)^n has to be in there somewhere, too

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I can't see a way to write an expression for the \(n\)th term without using imaginary numbers. That alternating pattern is tricky ... something like \[\large a_n=\frac{|i^n|}{n}\]

OpenStudy (anonymous):

I thought one might have to use imaginary numbers, since -1^n isn't enough

OpenStudy (anonymous):

I suppose there's nothing stopping you from writing \(a_n\) as a piecewise function that considers even and odd \(n\) separately.

OpenStudy (anonymous):

this is the answer wolframalpha gave: a_n = ((1/2+i/2) i^n (-i+(-1)^n))/n

OpenStudy (anonymous):

I think the formula I suggested is simpler... No telling what WA did in its calculations. As for the sum part of the question, you can try comparing to the series \(\displaystyle \sum\dfrac{(-1)^n}{n}\) if it's convergent, or maybe \(\displaystyle \sum\dfrac{1}{n}\) if it diverges.

OpenStudy (anonymous):

I eventually had to give up and just have the answer

OpenStudy (anonymous):

I tested what WA gave with n=10 -> n = 15, and it holds true. Putting that eqation in to try to look for simpler ones gave me ... i^n * e^(i*pi*n)

OpenStudy (anonymous):

WA says it converges to 0, compared to (-1)^n/n, which converges to -log2

OpenStudy (anonymous):

You mean the summation, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Strange... If you use \(\dfrac{|i^n|}{n}\), the series diverges, but it gives the same sequence.

OpenStudy (anonymous):

For the above equation, only positive fractions are outputs...

OpenStudy (anonymous):

Oh right, what was I thinking... The absolute value gets rid of the negatives. I meant the real part of \(i^n\).

OpenStudy (anonymous):

Wait, that's not right either... We'll need an expression for the sign of \(i^n\).

OpenStudy (anonymous):

The one you gave gives: { i, -1/2, -i/3, 1/4, i/5, .... }

OpenStudy (anonymous):

Well, at any rate you have a closed form for \(a_n\) that does work thanks to WA. How one would get is beyond me :/

OpenStudy (anonymous):

thanks for the help and input

OpenStudy (anonymous):

funny to see a math genius like you being boggled.

OpenStudy (anonymous):

Hardly a genius ;)

OpenStudy (anonymous):

by the way, for kicks, I looked at what WA gave for alternating fractions for every 3 terms...and it uses sine, i, and pi :P

OpenStudy (anonymous):

*actually, no i. maybe that one would be easier, I don't know. Thanks again.

OpenStudy (anonymous):

No problem